### v1

# A Geometric Area Bound for Information Transfer Through Semiclassical Traversable Wormholes

Abstract
We prove a rigorous geometric bound on the transmission of quantum information through semiclassical traversable wormholes. We initially prove an area theorem showing that the minimal throat surface of a traversable wormhole cannot increase using Raychaudhuri’s equation together with the null energy condition for the infalling matter. Then, the max-flow in the bit-thread picture, which corresponds to the initial minimal throat area, is shown to set the upper bound on information transfer. We also discuss two glued HaPPY codes as a toy model for a wormhole.
Introduction.— The interplay between entanglement and geometry has been a rapidly developing area of research in modern theoretical physics. One of the most concrete physical realizations of this relationship is the correspondence between the entanglement entropy of a boundary region and the area of the bulk minimal surface homologous to the boundary region, as prescribed by Ryu–Takayanagi and Hubeny–Rangamani–Takayanagi  [1, 2]. The eternal AdS-Schwarzschild black hole provides a canonical example of this correspondence, and is dual to the thermofield double state for two entangled quantum systems [3]. This duality is one of the most natural realizations of the broader “ER = EPR” idea [7], relating quantum entanglement to spacetime connectedness.
Classically, these wormholes are non-traversable; no signal entering one exterior can reach the other. This is because of the energy conditions imposed by general relativity. However, wormholes in general relativity can be made traversable once negative energy is introduced [4]. An eternal two-sided AdS-Schwarzschild black hole can be rendered traversable given a coupling between the two boundaries is introduced as it injects negative energy, violating the averaged null energy condition (ANEC), as shown by Gao, Jafferis, and Wall (GJW) [6].
The possibility of traversability raises a natural and broadly relevant question: How much information can be sent through such a wormhole at a time? Previous analyses have addressed this question in limited cases, often relying on nearly two-dimensional geometries or on perturbative backreaction arguments showing that the wormhole is closed after some information is thrown in [6, 18, 8, 9]. Despite this progress, a general and geometric principle has remained unknown.
In this Letter, we provide such a principle. We prove that for any semiclassical wormhole made traversable by a Gao-Jafferis-Wall deformation, the area of the wormhole throat, a smooth codimension-$2$ minimal surface, cannot be increased by any infalling matter satisfying the null energy condition (NEC).
By combining this geometric monotonicity with the max-flow/min-cut theorem in the bit threads formulation of holographic entropy [10], we derive a sharp upper bound for information transmission: the maximum number of independent quantum degrees of freedom that can be transmitted during the traversable window is exactly $A_{\min}/4G_{N}$. We further illustrate this capacity bound in a holographic tensor-network model built from entangled HaPPY codes applying the max-flow min-cut theorem. Our results provide a new area theorem for minimal wormhole throat surfaces and thus, give a bound on information transfer. Geometry of the wormhole throat and an area theorem.— We work in a smooth, time-oriented, globally hyperbolic spacetime $(M,g_{ab})$ of dimension $d\geq 3$ with two asymptotically AdS boundary components $\mathcal{I}_{L}$ and $\mathcal{I}_{R}$. The metric satisfies Einstein’s equations
$R_{ab}-\frac{1}{2}Rg_{ab}+\Lambda g_{ab}=8\pi G\,T_{ab},$ (1)
where $T_{ab}$ is the stress-energy tensor. The spacetime is assumed to contain a Gao-Jafferis-Wall (GJW) type traversable wormhole. The two boundaries are causally connected through the bulk due to a prior ANEC-violating deformation. After this deformation is switched off, we consider additional infalling matter that transfers information between the boundaries. This additional matter is taken to satisfy the null energy condition (NEC) along the null geodesics that carry the signals.
Let $\Sigma$ be a spacelike Cauchy slice intersecting the wormhole, and let $S_{0}\subset\Sigma$ be a smooth, compact, spacelike codimension-$2$ surface homologous to, say, the right boundary $\mathcal{I}_{R}$. $S_{0}$ is a cross-section of the wormhole throat. This geometry is depicted in Fig. 1(a). At each point of $S_{0}$, the normal bundle is $2$-dimensional and Lorentzian. We choose two future-directed null normals $k^{a}$ and $\ell^{a}$ that satisfy
$k^{a}\ell_{a}=-1.$ (2)
Figure 1: (Color online.)(a) Spatial Cauchy slice ($\Sigma$) with a wormhole whose minimal cross-section is the throat $S_{0}$, connecting the two asymptotic regions $\mathcal{I}_{L}$ and $\mathcal{I}_{R}$. (b) Null congruence on $\mathcal{N}$ with basis vectors $\ell^{a}$ and $k^{a}$.
The induced metric on $S_{0}$ is then
$h_{ab}=g_{ab}+k_{a}\ell_{b}+\ell_{a}k_{b}.$ (3)
The null expansion along $k^{a}$ is defined by
$\theta_{(k)}=k^{a}\nabla_{a}\!\left(\ln[Area(S_{0})]\right),$ (4)
and is deeply related to the change of the area element $\sqrt{h}$, where $h$ is the determinant of the metric $h_{ab}$, along the null congruence generated by $k^{a}$:
$\frac{d}{d\lambda}\ln\sqrt{h}=\theta_{(k)},\qquad\frac{d\sqrt{h}}{d\lambda}=\theta_{(k)}\sqrt{h},$ (5)
where $\lambda$ is an affine parameter along the null geodesics.
$k^{a}$ is chosen to point in the direction of traversal, e.g., from $\mathcal{I}_{L}$ to $\mathcal{I}_{R}$ through the bulk. We impose the following geometric condition on the throat cross-section:
$\theta_{(k)}\big|_{S_{0}}\leq 0.$ (6)
Physically, this reflects that $S_{0}$ is a bottleneck for signals propagating along $k^{a}$: the cross-sectional area neither expands nor contracts at the throat in that null direction. This condition is weaker than requiring $S_{0}$ to be an ext remal surface in the full spacetime, as it does not require $\theta^{(k)}=0$ or $\theta^{(\ell)}=0$ individually.
Let $\mathcal{N}$ be the null hypersurface generated by future-directed, affinely parameterized null geodesics $\gamma(\lambda)$ starting on $S_{0}$ with tangent $k^{a}$:
$\mathcal{N}=\big\{\gamma(\lambda)\,\big|\,\gamma(0)\in S_{0},\ \dot{\gamma}^{a}(0)=k^{a},\ \lambda\geq 0\big\}.$ (7)
By construction, $\mathcal{N}$ is hypersurface-orthogonal, so the vorticity of the congruence vanishes. For each $\lambda>0$ before caustics, we define a cross-section $S(\lambda)$ as a spacelike codimension-$2$ slice of $\mathcal{N}$ that intersects each generator only once. An illustration of $\mathcal{N}$ is given in Fig. 1(b). The area of $S(\lambda)$ is denoted
$A(\lambda):=\mathrm{Area}\big(S(\lambda)\big).$ (8)
The sign-preserving structure of the null expansion plays a crucial role in the proof of the first theorem to be proven.
Lemma 1.
Lemma. Let $\mathcal{N}$ be a null hypersurface generated by affinely parameterized null geodesics with tangent $k^{a}$, and let $\theta(\lambda)$ be the expansion along $k^{a}$ on a given generator. Assume that along $\mathcal{N}$ the infalling matter satisfies the null energy condition,
$T_{ab}k^{a}k^{b}\geq 0,$ (9)
and that the congruence has vanishing twist, as is the case for a null hypersurface. If $\theta(\lambda_{0})\leq 0$ at some $\lambda_{0}$ on a generator, then
$\theta(\lambda)\leq 0\quad\text{for all }\lambda\geq\lambda_{0},$ (10)
up to the first conjugate point.
Proof.
For a hypersurface-orthogonal null congruence, the Raychaudhuri equation is given by
$\frac{d\theta}{d\lambda}=-\frac{1}{2}\theta^{2}-\sigma_{ab}\sigma^{ab}-R_{ab}k^{a}k^{b},$ (11)
where $\sigma_{ab}$ is the shear tensor. Contracting Eq. (1) with $k^{a}k^{b}$ and using $g_{ab}k^{a}k^{b}=0$ yields
$R_{ab}k^{a}k^{b}=8\pi G\,T_{ab}k^{a}k^{b}.$ (12)
Substituting this into Eq.  (11), using $\sigma_{ab}\sigma^{ab}\geq 0$, and the NEC (9), we obtain
$\frac{d\theta}{d\lambda}\leq-\frac{1}{2}\theta^{2}.$ (13)
Assume that $\theta(\lambda_{0})\leq 0$ at some $\lambda_{0}$. Also, suppose that there exists $\lambda_{1}>\lambda_{0}$ with $\theta(\lambda_{1})>0$. By continuity of $\theta(\lambda)$, there is a first $\lambda_{*}\in(\lambda_{0},\lambda_{1}]$ such that
$\theta(\lambda_{*})=0,\qquad\theta(\lambda)<0\ \text{for}\ \lambda_{0}\leq\lambda<\lambda_{*}.$ (14)
Evaluating Eq. (11) at $\lambda_{*}$, we have
$\left.\frac{d\theta}{d\lambda}\right|_{\lambda_{*}}=-\sigma_{ab}\sigma^{ab}-8\pi G\,T_{ab}k^{a}k^{b}\leq 0,$ (15)
since both terms are non-negative. However, for $\theta(\lambda)$ to transition from negative values for $\lambda<\lambda_{*}$ to positive values for some $\lambda>\lambda_{*}$, the derivative at the first zero must satisfy $\left.d\theta/d\lambda\right|_{\lambda_{*}}>0$. This is a contradiction. Therefore $\theta(\lambda)\leq 0$ for all $\lambda\geq\lambda_{0}$ up to the first conjugate point. ∎
We can now state and prove the main geometric result.
Theorem 1.
Let $S_{0}$ be a wormhole throat cross-section in a GJW-type traversable AdS two-sided spacetime, as described previously, and let $\mathcal{N}$ be the null hypersurface generated from $S_{0}$ by future-directed null geodesics orthogonal to $S_{0}$ with tangent $k^{a}$ pointing along the direction of traversal. Assume that
  1. 1.
    The additional infalling matter used to send signals through the wormhole satisfies the null energy condition along $\mathcal{N}$, $T_{ab}k^{a}k^{b}\geq 0$.
  2. 2.
    The throat condition (6) holds on $S_{0}$, $\theta_{(k)}\big|_{S_{0}}\leq 0$.
  3. 3.
    There are no conjugate points along $\mathcal{N}$, so that each null generator intersects each cross-section $S(\lambda)$ exactly once.
Then the area $A(\lambda)$ of the cross-sections $S(\lambda)\subset\mathcal{N}$ is non-increasing toward the future:
$\frac{dA}{d\lambda}\leq 0,$ (16)
with equality being true for an open interval in $\lambda$ only if $\theta=0$, $\sigma_{ab}=0$, and $T_{ab}k^{a}k^{b}=0$ along $\mathcal{N}$ in that interval.
Proof.
By assumption, on each generator of $\mathcal{N}$ the condition $\theta(0)\leq 0$ at $\lambda_{0}=0$, the location of $S_{0}$, is satisfied. The NEC holds along $\mathcal{N}$, the congruence is twist-free, and any conjugates in the area of interest are excluded. By Lemma 1, we have
$\theta(\lambda)\leq 0\quad\text{for all }\lambda\geq 0$ (17)
on each generator, up to a conjugate point.
Let $\sqrt{h}(\lambda,x)$ be the area element on $S(\lambda)$, where $x$ stands for the generators. Equation (5) implies that
$\frac{d}{d\lambda}\sqrt{h}(\lambda,x)=\theta(\lambda,x)\,\sqrt{h}(\lambda,x).$ (18)
Since $\theta(\lambda,x)\leq 0$ pointwise and $\sqrt{h}(\lambda,x)\geq 0$, it follows that
$\frac{d}{d\lambda}\sqrt{h}(\lambda,x)\leq 0\quad\text{for all }x,\lambda\geq 0.$ (19)
The area of $S(\lambda)$ is given by
$A(\lambda)=\int_{S(\lambda)}\sqrt{h}(\lambda,x)\,d^{d-2}x.$ (20)
Differentiating under the integral sign, we obtain
$\frac{dA}{d\lambda}=\int_{S(\lambda)}\theta(\lambda,x)\,\sqrt{h}(\lambda,x)\,d^{d-2}x\leq 0,$ (21)
since the integrand is everywhere non-positive. This proves the claimed monotonicity.
If $dA/d\lambda=0$ on some open interval in $\lambda$, then the integral of the non-positive integrand must vanish, which implies $\theta(\lambda,x)=0$ for all $x$ on that interval. Substituting $\theta=0$ into the Raychaudhuri equation (11) yields
$0=-\sigma_{ab}\sigma^{ab}-8\pi G\,T_{ab}k^{a}k^{b},$ (22)
so $\sigma_{ab}=0$ and $T_{ab}k^{a}k^{b}=0$ along $\mathcal{N}$ in that interval. Thus, the area can remain strictly constant only when the congruence is shear-free, and there is no null energy flux along the generators. ∎
Theorem 1 is a classical area law for null congruences originating from a traversable wormhole throat in the post-ANEC violation regime. In particular, it implies that any NEC-satisfying signal sent through the wormhole cannot increase the cross-sectional area of the throat along the direction of traversal
Bounds on information transfer and bit threads. — Let $\Sigma$ be a Cauchy slice intersecting the traversable wormhole region, and equip it with the induced Riemannian metric $h_{ab}$. We regard $\Sigma$ as a spatial manifold with two asymptotic boundaries corresponding to $\mathcal{I}_{L}$ and $\mathcal{I}_{R}$. Any causal curve connecting $\mathcal{I}_{L}$ to $\mathcal{I}_{R}$ must intersect some codimension-$1$ hypersurface on $\Sigma$ that separates the two sides, and in particular must intersect any codimension-$2$ surface homologous to a boundary that lies on such a separation.
By construction, $\gamma_{\min}$ is a codimension-$2$ surface that separates the left and right sides and has minimal area among extremal surfaces of this type. Theorem 1 implies that, within the causal development of $\Sigma$, no null deformation can increase the area of $\gamma_{\min}$ beyond the area $Area(\gamma_{min})\equiv{A_{min}}$. In particular, the wormhole throat is a cross-section of area at least $A_{\min}$, and $\gamma_{\min}$ is realized as the minimal extremal surface.
On the Riemannian manifold $(\Sigma,h_{ab})$, Freedman and Headrick introduced divergenceless vector fields, bit threads, $v^{a}$ whose norms are bounded by [10]
$\nabla_{a}v^{a}=0,\qquad|v|\leq\frac{1}{4G_{N}},$ (23)
and proved a max-flow/min-cut theorem stating that the maximum flux of such a field from $\mathcal{I}_{L}$ to $\mathcal{I}_{R}$ equals the minimal area of any separating surface divided by $4G_{N}$, equivalently, the entanglement entropy between the two boundaries,
$S_{EE}=\max_{v}\Phi[v]=\frac{A_{\min}}{4G_{N}},$ (24)
where the flux $\Phi[v]$ is the integral of $v^{a}$ over a boundary region. Bit threads thus represent the entanglement between different points on boundaries.
In the holographic context, such flows can be interpreted as ensembles of bit threads whose density is bounded by (23); the flux (24) then counts the maximal number of distinguishable threads that can emanate from one boundary to the other.
To turn this geometric statement into an information-theoretic bound, we consider any information transfer protocol that uses the traversable wormhole once during the interval in which signals can propagate from $\mathcal{I}_{L}$ to $\mathcal{I}_{R}$. The protocol transmits $Q$ independent quantum degrees of freedom if there exists a set of $Q$ mutually orthogonal input states on $\mathcal{I}_{L}$ whose corresponding output states on $\mathcal{I}_{R}$ remain mutually orthogonal and thus perfectly distinguishable by some measurement on $\mathcal{I}_{R}$ alone.
Each such distinguishable output must be supported on a distinct pattern of bit threads, because two states that occupy exactly the same set of threads with the same local densities along the entire traversable region cannot be reliably distinguished at $\mathcal{I}_{R}$, as their entanglement structure with the left boundary is the same. As the bit threads across the boundaries represent shared entangled qubits, there is a natural upper bound on the information that each thread can carry. A single external qubit can be teleported from $I_{L}$ using two bits of classical information and one shared Bell pair; hence each bit thread can transmit at most one qubit. Therefore $S_{\mathrm{EE}}=A_{\min}/4G_{N}$ non-intersecting bit threads can teleport up to $S_{\mathrm{EE}}$ qubits. This argument sets the fundamental upper bound. One should note that quantum teleportation is often argued to be dual to information transfer through a traversable wormhole, supported by ideas from AdS/CFT [18].
Let there be two asymptotic regions $L$ and $R$, which stand for black hole boundaries on the left and right, respectively, in addition to some infalling matter M. The conditional–mutual–information flow v(R:L|M)v(R\!:\!L\,|\,M) provides another way to state the geometric bound. This divergenceless bit threads flow is supported only in the tube consisting of entanglement wedges of $r(RLM)\setminus r(M)$ connecting $R$ and $L$, and its flux is bounded by the smallest area surface separating the two sides within this tube, which is given by the conditional mutual information
$I(R\!:L\,|\,M)=\frac{A^{\prime}_{min}}{4G_{N}},$ (25)
where $A^{\prime}_{min}$ is the area of the minimal surface after the infall of the additional matter, which we have proved cannot be larger than $A_{min}$. In this context, it is reminiscent that the conditional mutual information is more fundamental than the entanglement entropy in terms of a quantum information theoretic bound, and the entanglement entropy is the special case when the infalling matter is not correlated with either of the boundaries. Thus, bit threads fit naturally into the picture, and their flux provides an upper bound on information transfer.
A protocol capable of transmitting $Q$ mutually orthogonal states requires access to at least $Q$ mutually distinguishable thread flux patterns. This operational condition directly motivates the identification of the channel capacity with the maximal bit threads flux and leads to the geometric upper bound established in Theorem 2.
Theorem 2.
Let a two-sided asymptotically AdSd+1 wormhole be rendered traversable by a Gao-Jafferis-Wall deformation in a semiclassical spacetime. Let $\gamma_{\min}$ be a smooth, compact, minimal cross-section. Then the maximum number $Q_{\max}$ of independent quantum degrees of freedom that can be transmitted from one boundary to the other during the traversable interval satisfies
$Q_{\max}=\frac{A_{\min}}{4G_{N}}.$ (26)
Proof. The argument above shows that any protocol capable of transmitting $Q$ mutually orthogonal input states must realize at least $Q$ units of flux through $\gamma_{\min}$ when described in terms of a divergenceless flow $v^{a}$ obeying (23). By the max-flow/min-cut theorem (24), such a flow must satisfy $\Phi[v]\leq A_{\min}/4G_{N}$. Hence $Q\leq A_{\min}/4G_{N}$. Taking the supremum over all protocols gives $Q_{\max}\leq A_{\min}/4G_{N}$. On the other hand, standard quantum teleportation arguments show that one can saturate this inequality asymptotically by sending one quantum degree of freedom per bit threads, so that $Q_{\max}=A_{\min}/4G_{N}$ in the idealized limit. ∎
Theorem 2 essentially states that a traversable wormhole cannot serve as a quantum channel with capacity exceeding one quantum degree of freedom per Planck area of its minimal extremal cross-section. This is the natural transmission analogue of the Bekenstein–Hawking bound on information storage [19, 20, 21].
Holographic code toy model. — The geometric capacity bound has a simple analogue in a tensor-network toy model. Two HaPPY codes can be glued along their bulk legs to form a discrete two-sided wormhole, where the minimal cut $\Gamma_{\min}$ plays the role of the throat. The number of bonds $n_{\min}$ cut by $\Gamma_{\min}$ is the discrete counterpart of the minimal area, and the max-flow/min-cut theorem implies that at most $n_{\min}$ independent qubits can be transferred from one boundary to the other as the capacity of each leg connecting the two black holes is $S_{BH}/n_{min}$. Thus, the toy model reproduces the continuum capacity bound, the wormhole transmits no more independent degrees of freedom than the number of bit threads that can fit across its minimal cross-section.
Discussion. — We have shown that traversable wormholes obey a sharp geometric bound on information transfer, controlled solely by the minimal extremal cross-section, the wormhole throat. Our findings provide a compact geometric framework for understanding the limits of information flow in semiclassical gravity.
Once the GJW deformation is switched off, the throat area cannot increase under any additional matter satisfying the NEC. This monotonicity constrains the maximal information flux independently of the dynamical mechanism that limits the duration of traversability. Backreaction effects may reduce the effective time available for transmission, but they cannot allow capacities exceeding those permitted by the minimal cross-section. Thus, dynamical closure and the geometric constraint are fully compatible, with the latter providing the more fundamental limitation.
It is natural to ask how the present analysis extends beyond the classical setting. Our results apply to classical extremal surfaces and rely on the Raychaudhuri equation for null congruences. In a quantum-corrected regime, extremal surfaces are replaced by quantum extremal surfaces (QES), for which the generalized entropy, $S_{gen}$,
$S_{gen}=\frac{Area(\gamma)}{4G_{N}}+S_{bulk}(\gamma)$ (27)
is extremized  [22]. The monotonicity properties of $S_{gen}$ involve additional contributions from bulk entanglement $S_{bulk}$, and a modified Raychaudhuri-type framework would be required to establish an analogous area theorem. Developing such a “quantum” version of the capacity bound is an interesting open problem that we leave for future work.
Because the throat capacity matches the Page-time interior entropy, our result suggests a geometric mechanism underlying the Hayden–Preskill protocol [12]. A traversable wormhole can be used to support a quantum channel capable of transmitting all interior black hole degrees of freedom that remain entangled with the Hawking radiation, offering a way to extract all the information trapped behind the horizon. It is interesting to note that the bound on information transfer is precisely what is needed to prevent information loss. It would be fruitful to explore this rather speculative topic on mathematical grounds.
The results presented here indicate that traversable wormholes possess a universal geometric bottleneck for information transfer, determined solely by the minimal extremal cross-section. The equivalence between this geometric constraint and the entanglement-theoretic max-flow strongly suggests a deeper structural relation between semiclassical geometry and quantum information. Extending this correspondence to quantum extremal surfaces and generalized entropy, and the mathematical study of the applications to the black hole information problem, remain intriguing directions for future work.

### v2

@@
-A Geometric Area Bound for Information Transfer Through Semiclassical Traversable Wormholes
+An Area Theorem for Traversable Wormhole Throats
 Abstract
-We prove a rigorous geometric bound on the transmission of quantum information through semiclassical traversable wormholes.
-We initially prove an area theorem showing that the minimal throat surface of a traversable wormhole cannot increase using Raychaudhuri’s equation together with the null energy condition for the infalling matter.
-Then, the max-flow in the bit-thread picture, which corresponds to the initial minimal throat area, is shown to set the upper bound on information transfer.
-We also discuss two glued HaPPY codes as a toy model for a wormhole.
+We prove a new area theorem in general relativity showing that, after a Gao-Jafferis-Wall deformation is switched off, the area of a wormhole throat cross-section cannot increase under infalling matter obeying the null energy condition.
+As a consequence, the minimal throat area sets a sharp upper bound on information transfer.
+Using the max-flow/min-cut formulation of bit threads, we show that the number of independent quantum degrees of freedom that can be transmitted through a semiclassical traversable wormhole is bounded by the throat area.
+We further illustrate the bound with a holographic tensor-network toy model built from two glued HaPPY codes.
 Introduction.— The interplay between entanglement and geometry has been a rapidly developing area of research in modern theoretical physics.
-One of the most concrete physical realizations of this relationship is the correspondence between the entanglement entropy of a boundary region and the area of the bulk minimal surface homologous to the boundary region, as prescribed by Ryu–Takayanagi and Hubeny–Rangamani–Takayanagi  [1, 2].
-The eternal AdS-Schwarzschild black hole provides a canonical example of this correspondence, and is dual to the thermofield double state for two entangled quantum systems [3].
+One of the most concrete physical realizations of this relationship is the correspondence between the entanglement entropy of a boundary region and the area of the minimal bulk surface homologous to the boundary region, as prescribed by Ryu–Takayanagi and Hubeny–Rangamani–Takayanagi  [1, 2].
+The eternal AdS-Schwarzschild black hole provides a canonical example of this correspondence and is dual to the thermofield double state for two entangled quantum systems [3].
 This duality is one of the most natural realizations of the broader “ER = EPR” idea [4], relating quantum entanglement to spacetime connectedness.
@@
 This is because of the energy conditions imposed by general relativity.
-However, wormholes in general relativity can be made traversable once negative energy is introduced [4].
-An eternal two-sided AdS-Schwarzschild black hole can be rendered traversable given a coupling between the two boundaries is introduced as it injects negative energy, violating the averaged null energy condition (ANEC), as shown by Gao, Jafferis, and Wall (GJW) [6].
+However, wormholes in general relativity can be traversable once negative energy is introduced [5].
+An eternal two-sided AdS-Schwarzschild black hole can be rendered traversable by introducing a coupling between the two boundaries that injects negative energy and violates the averaged null energy condition (ANEC), as shown by Gao, Jafferis, and Wall (GJW) [6].
 The possibility of traversability raises a natural and broadly relevant question: How much information can be sent through such a wormhole at a time?
 Previous analyses have addressed this question in limited cases, often relying on nearly two-dimensional geometries or on perturbative backreaction arguments showing that the wormhole is closed after some information is thrown in [6, 9, 7, 8].
-Despite this progress, a general and geometric principle has remained unknown.
-In this Letter, we provide such a principle.
-We prove that for any semiclassical wormhole made traversable by a Gao-Jafferis-Wall deformation, the area of the wormhole throat, a smooth codimension-$2$ minimal surface, cannot be increased by any infalling matter satisfying the null energy condition (NEC).
-By combining this geometric monotonicity with the max-flow/min-cut theorem in the bit threads formulation of holographic entropy [10], we derive a sharp upper bound for information transmission: the maximum number of independent quantum degrees of freedom that can be transmitted during the traversable window is exactly $A_{\min}/4G_{N}$.
-We further illustrate this capacity bound in a holographic tensor-network model built from entangled HaPPY codes applying the max-flow min-cut theorem.
-Our results provide a new area theorem for minimal wormhole throat surfaces and thus, give a bound on information transfer.
-Geometry of the wormhole throat and an area theorem.— We work in a smooth, time-oriented, globally hyperbolic spacetime $(M,g_{ab})$ of dimension $d\geq 3$ with two asymptotically AdS boundary components $\mathcal{I}_{L}$ and $\mathcal{I}_{R}$.
+No general area theorem for traversable wormhole throats or universal bound on their communication capacity is known.
+In this Letter, we provide such a theorem and bound.
+We prove that after the GJW deformation is switched off, the area of the wormhole throat, a smooth codimension-$2$ cross-section, cannot be increased by any infalling matter satisfying the null energy condition (NEC).
+By combining this geometric monotonicity with the max-flow/min-cut theorem in the bit thread formulation of holographic entropy [10], we derive a sharp upper bound for information transmission: the maximum number of independent quantum degrees of freedom that can be transmitted during the traversable window is at most $A_{\min}/4G_{N}$.
+We further illustrate this capacity bound in a holographic tensor-network model built from two entangled black holes in two glued HaPPY codes applying the max-flow/min-cut theorem.
+Our results provide a new area theorem for traversable wormhole throat surfaces, and thus give a bound on information transfer.
+Although our analysis involves null congruences and an area–entropy relation, it is conceptually distinct from the Bousso covariant entropy bound [11], which constrains entropy on light-sheets rather than the geometry or capacity of a wormhole throat.
+Unlike Hawking’s area theorem [12], which applies to event horizons, our result constrains arbitrary wormhole throat cross-sections that act as geometric bottlenecks for information transfer.
+An area theorem.— We work in a smooth, time-oriented spacetime $(M,g_{ab})$ of dimension $D\geq 3$ with two asymptotically AdS boundary components $\mathcal{I}_{L}$ and $\mathcal{I}_{R}$.
 The metric satisfies Einstein’s equations
@@
 This additional matter is taken to satisfy the null energy condition (NEC) along the null geodesics that carry the signals.
-Let $\Sigma$ be a spacelike Cauchy slice intersecting the wormhole, and let $S_{0}\subset\Sigma$ be a smooth, compact, spacelike codimension-$2$ surface homologous to, say, the right boundary $\mathcal{I}_{R}$.
+Let $\Sigma$ be a spacelike hypersurface intersecting the wormhole, and let $S_{0}\subset\Sigma$ be a smooth, compact, spacelike codimension-$2$ surface homologous to, say, the right boundary $\mathcal{I}_{R}$.
 $S_{0}$ is a cross-section of the wormhole throat.
@@
 1(a).
-At each point of $S_{0}$, the normal bundle is $2$-dimensional and Lorentzian.
+At each point of $S_{0}$, the normal bundle is 22–dimensional and Lorentzian.
 We choose two future-directed null normals $k^{a}$ and $\ell^{a}$ that satisfy
 $k^{a}\ell_{a}=-1.$ (2)
-Figure 1: (Color online.)(a) Spatial Cauchy slice ($\Sigma$) with a wormhole whose minimal cross-section is the throat $S_{0}$, connecting the two asymptotic regions $\mathcal{I}_{L}$ and $\mathcal{I}_{R}$.
+Figure 1: (a) Spacelike hypersurface ($\Sigma$) intersecting a wormhole, whose minimal cross-section is the throat $S_{0}$, connecting the two asymptotic regions $\mathcal{I}_{L}$ and $\mathcal{I}_{R}$.
 (b) Null congruence on $\mathcal{N}$ with basis vectors $\ell^{a}$ and $k^{a}$.
@@
 The null expansion along $k^{a}$ is defined by
-$\theta_{(k)}=k^{a}\nabla_{a}\!\left(\ln[Area(S_{0})]\right),$ (4)
+$\theta_{(k)}=h^{ab}\nabla_{a}k_{b},$ (4)
 and is deeply related to the change of the area element $\sqrt{h}$, where $h$ is the determinant of the metric $h_{ab}$, along the null congruence generated by $k^{a}$:
@@
 $\theta_{(k)}\big|_{S_{0}}\leq 0.$ (6)
-Physically, this reflects that $S_{0}$ is a bottleneck for signals propagating along $k^{a}$: the cross-sectional area neither expands nor contracts at the throat in that null direction.
-This condition is weaker than requiring $S_{0}$ to be an ext remal surface in the full spacetime, as it does not require $\theta^{(k)}=0$ or $\theta^{(\ell)}=0$ individually.
+Physically, this reflects that $S_{0}$ is a bottleneck for signals propagating along $k^{a}$: the cross-sectional area is non-increasing at the throat in that null direction.
+This condition is weaker than requiring $S_{0}$ to be an extremal surface in the full spacetime, as it does not require $\theta_{(k)}=0$ or $\theta_{(\ell)}=0$ individually.
 Let $\mathcal{N}$ be the null hypersurface generated by future-directed, affinely parameterized null geodesics $\gamma(\lambda)$ starting on $S_{0}$ with tangent $k^{a}$:
@@
 The area of $S(\lambda)$ is denoted
-$A(\lambda):=\mathrm{Area}\big(S(\lambda)\big).$ (8)
+$A(\lambda):=\text{Area}\big(S(\lambda)\big).$ (8)
 The sign-preserving structure of the null expansion plays a crucial role in the proof of the first theorem to be proven.
 Lemma 1.
-Lemma.
 Let $\mathcal{N}$ be a null hypersurface generated by affinely parameterized null geodesics with tangent $k^{a}$, and let $\theta(\lambda)$ be the expansion along $k^{a}$ on a given generator.
@@
 For a hypersurface-orthogonal null congruence, the Raychaudhuri equation is given by
-$\frac{d\theta}{d\lambda}=-\frac{1}{2}\theta^{2}-\sigma_{ab}\sigma^{ab}-R_{ab}k^{a}k^{b},$ (11)
+$\frac{d\theta}{d\lambda}=-\frac{1}{D-2}\theta^{2}-\sigma_{ab}\sigma^{ab}-R_{ab}k^{a}k^{b},$ (11)
 where $\sigma_{ab}$ is the shear tensor. Contracting Eq. (1) with $k^{a}k^{b}$ and using $g_{ab}k^{a}k^{b}=0$ yields
 $R_{ab}k^{a}k^{b}=8\pi G\,T_{ab}k^{a}k^{b}.$ (12)
-Substituting this into Eq.  (11), using $\sigma_{ab}\sigma^{ab}\geq 0$, and the NEC (9), we obtain
-$\frac{d\theta}{d\lambda}\leq-\frac{1}{2}\theta^{2}.$ (13)
+Substituting this into Eq. (11) and using $\sigma_{ab}\sigma^{ab}\geq 0$ together with the NEC (9), the Raychaudhuri equation reduces to
+$\frac{d\theta}{d\lambda}\leq-\frac{1}{D-2}\theta^{2}.$ (13)
 Assume that $\theta(\lambda_{0})\leq 0$ at some $\lambda_{0}$.
@@
 By continuity of $\theta(\lambda)$, there is a first $\lambda_{*}\in(\lambda_{0},\lambda_{1}]$ such that
-$\theta(\lambda_{*})=0,\qquad\theta(\lambda)<0\ \text{for}\ \lambda_{0}\leq\lambda<\lambda_{*}.$ (14)
+$\theta(\lambda_{*})=0,\qquad\theta(\lambda)<0\ \text{for}\ \lambda_{0}<\lambda<\lambda_{*}.$ (14)
 Evaluating Eq. (11) at $\lambda_{*}$, we have
@@
 Theorem 1.
-Let $S_{0}$ be a wormhole throat cross-section in a GJW-type traversable AdS two-sided spacetime, as described previously, and let $\mathcal{N}$ be the null hypersurface generated from $S_{0}$ by future-directed null geodesics orthogonal to $S_{0}$ with tangent $k^{a}$ pointing along the direction of traversal.
+Let $S_{0}$ be a wormhole throat cross-section in a GJW-type traversable AdS wormhole spacetime, as described previously, and let $\mathcal{N}$ be the null hypersurface generated from $S_{0}$ by future-directed null geodesics orthogonal to $S_{0}$ with tangent $k^{a}$ pointing along the direction of traversal.
 Assume that
@@
     The throat condition (6) holds on $S_{0}$, $\theta_{(k)}\big|_{S_{0}}\leq 0$.
-  3. 3.
-    There are no conjugate points along $\mathcal{N}$, so that each null generator intersects each cross-section $S(\lambda)$ exactly once.
-Then the area $A(\lambda)$ of the cross-sections $S(\lambda)\subset\mathcal{N}$ is non-increasing toward the future:
+Then the area $A(\lambda)$ of the cross-sections $S(\lambda)\subset\mathcal{N}$ is non-increasing toward the future up to the first conjugate point along $\mathcal{N}$, so that each null generator intersects each cross-section $S(\lambda)$ exactly once:
 $\frac{dA}{d\lambda}\leq 0,$ (16)
@@
 Bounds on information transfer and bit threads.
-— Let $\Sigma$ be a Cauchy slice intersecting the traversable wormhole region, and equip it with the induced Riemannian metric $h_{ab}$.
+— Let $\Sigma$ be a spacelike hypersurface intersecting the traversable wormhole region, and equip it with the induced Riemannian metric $h_{ab}$.
 We regard $\Sigma$ as a spatial manifold with two asymptotic boundaries corresponding to $\mathcal{I}_{L}$ and $\mathcal{I}_{R}$.
 Any causal curve connecting $\mathcal{I}_{L}$ to $\mathcal{I}_{R}$ must intersect some codimension-$1$ hypersurface on $\Sigma$ that separates the two sides, and in particular must intersect any codimension-$2$ surface homologous to a boundary that lies on such a separation.
-By construction, $\gamma_{\min}$ is a codimension-$2$ surface that separates the left and right sides and has minimal area among extremal surfaces of this type.
+By construction, $\gamma_{\min}$ is a codimension-$2$ surface that separates the left and right sides and has minimal area among codimension-2 surfaces on $\Sigma$ homologous to a boundary of this type.
 Theorem 1 implies that, within the causal development of $\Sigma$, no null deformation can increase the area of $\gamma_{\min}$ beyond the area $Area(\gamma_{min})\equiv{A_{min}}$.
-In particular, the wormhole throat is a cross-section of area at least $A_{\min}$, and $\gamma_{\min}$ is realized as the minimal extremal surface.
+In particular, the wormhole throat is a cross-section of area at least $A_{\min}$, and $\gamma_{\min}$ is realized as the minimal extremal surface satisfying $\gamma_{\min}=S_{0}$ on $\Sigma$.
 On the Riemannian manifold $(\Sigma,h_{ab})$, Freedman and Headrick introduced divergenceless vector fields, bit threads, $v^{a}$ whose norms are bounded by [10]
@@
 where the flux $\Phi[v]$ is the integral of $v^{a}$ over a boundary region. Bit threads thus represent the entanglement between different points on boundaries.
-In the holographic context, such flows can be interpreted as ensembles of bit threads whose density is bounded by (23); the flux (24) then counts the maximal number of distinguishable threads that can emanate from one boundary to the other.
+In the holographic context, such flows can be interpreted as ensembles of bit threads whose densities are bounded by (23); the flux (24) then counts the maximal number of distinguishable threads that can emanate from one boundary to the other.
 To turn this geometric statement into an information-theoretic bound, we consider any information transfer protocol that uses the traversable wormhole once during the interval in which signals can propagate from $\mathcal{I}_{L}$ to $\mathcal{I}_{R}$.
@@
 As the bit threads across the boundaries represent shared entangled qubits, there is a natural upper bound on the information that each thread can carry.
-A single external qubit can be teleported from $I_{L}$ using two bits of classical information and one shared Bell pair; hence each bit thread can transmit at most one qubit.
+A single external qubit can be teleported from $\mathcal{I}_{L}$ using two bits of classical information and one shared Bell pair; giving an operational upper bound of one qubit per thread.
 Therefore $S_{\mathrm{EE}}=A_{\min}/4G_{N}$ non-intersecting bit threads can teleport up to $S_{\mathrm{EE}}$ qubits.
@@
 One should note that quantum teleportation is often argued to be dual to information transfer through a traversable wormhole, supported by ideas from AdS/CFT [9].
-Let there be two asymptotic regions $L$ and $R$, which stand for black hole boundaries on the left and right, respectively, in addition to some infalling matter M.
-The conditional–mutual–information flow v(R:L|M)v(R\!:\!L\,|\,M) provides another way to state the geometric bound.
-This divergenceless bit threads flow is supported only in the tube consisting of entanglement wedges of $r(RLM)\setminus r(M)$ connecting $R$ and $L$, and its flux is bounded by the smallest area surface separating the two sides within this tube, which is given by the conditional mutual information
-$I(R\!:L\,|\,M)=\frac{A^{\prime}_{min}}{4G_{N}},$ (25)
-where $A^{\prime}_{min}$ is the area of the minimal surface after the infall of the additional matter, which we have proved cannot be larger than $A_{min}$.
-In this context, it is reminiscent that the conditional mutual information is more fundamental than the entanglement entropy in terms of a quantum information theoretic bound, and the entanglement entropy is the special case when the infalling matter is not correlated with either of the boundaries.
-Thus, bit threads fit naturally into the picture, and their flux provides an upper bound on information transfer.
 A protocol capable of transmitting $Q$ mutually orthogonal states requires access to at least $Q$ mutually distinguishable thread flux patterns.
@@
 Then the maximum number $Q_{\max}$ of independent quantum degrees of freedom that can be transmitted from one boundary to the other during the traversable interval satisfies
-$Q_{\max}=\frac{A_{\min}}{4G_{N}}.$ (26)
+$Q_{\max}\leq\frac{A_{\min}}{4G_{N}}.$ (25)
 Proof.
@@
 Taking the supremum over all protocols gives $Q_{\max}\leq A_{\min}/4G_{N}$.
-On the other hand, standard quantum teleportation arguments show that one can saturate this inequality asymptotically by sending one quantum degree of freedom per bit threads, so that $Q_{\max}=A_{\min}/4G_{N}$ in the idealized limit.
+On the other hand, standard quantum teleportation arguments show that one can saturate this inequality asymptotically by sending one quantum degree of freedom per bit thread, so that $Q_{\max}=A_{\min}/4G_{N}$ in the idealized limit.
 ∎
@@
 This is the natural transmission analogue of the Bekenstein–Hawking bound on information storage [13, 14, 15].
+From a quantum-information perspective, the capacity bound can be viewed as a statement about the conditional mutual information $I(\mathcal{I}_{R}:\mathcal{I}_{L}\mid M)$, where $M$ denotes the infalling matter system, reducing to the entanglement entropy when the infalling matter is uncorrelated with the black hole degrees of freedom, as naturally captured in the bit-thread formulation of holographic entanglement entropy [10].
 Holographic code toy model.
-— The geometric capacity bound has a simple analogue in a tensor-network toy model.
-Two HaPPY codes can be glued along their bulk legs to form a discrete two-sided wormhole, where the minimal cut $\Gamma_{\min}$ plays the role of the throat.
-The number of bonds $n_{\min}$ cut by $\Gamma_{\min}$ is the discrete counterpart of the minimal area, and the max-flow/min-cut theorem implies that at most $n_{\min}$ independent qubits can be transferred from one boundary to the other as the capacity of each leg connecting the two black holes is $S_{BH}/n_{min}$.
+— The geometric capacity bound has a simple analogue in a holographic tensor-network toy model, particularly the HaPPY code [16].
+Two black holes that are constructed by removing the central qubits in two HaPPY codes can be entangled along their bulk legs to form a discrete two-sided wormhole.
+Making use of the network structure of the HaPPY codes, the max-flow/min-cut theorem can be applied where the minimal cut $\Gamma_{\min}$ plays the role of the throat.
+The number of bonds $n_{\min}$ cut by $\Gamma_{\min}$ is the discrete counterpart of the minimal area, and the max-flow/min-cut theorem implies that at most $n_{\min}$ independent qubits can be transferred from one boundary to the other as the capacity of each leg connecting the two black holes one qubit.
 Thus, the toy model reproduces the continuum capacity bound, the wormhole transmits no more independent degrees of freedom than the number of bit threads that can fit across its minimal cross-section.
@@
 In a quantum-corrected regime, extremal surfaces are replaced by quantum extremal surfaces (QES), for which the generalized entropy, $S_{gen}$,
-$S_{gen}=\frac{Area(\gamma)}{4G_{N}}+S_{bulk}(\gamma)$ (27)
+$S_{gen}=\frac{\text{Area}(\gamma)}{4G_{N}}+S_{\text{bulk}}(\gamma)$ (26)
 is extremized  [17].
@@
 Developing such a “quantum” version of the capacity bound is an interesting open problem that we leave for future work.
-Because the throat capacity matches the Page-time interior entropy, our result suggests a geometric mechanism underlying the Hayden–Preskill protocol [12].
-A traversable wormhole can be used to support a quantum channel capable of transmitting all interior black hole degrees of freedom that remain entangled with the Hawking radiation, offering a way to extract all the information trapped behind the horizon.
-It is interesting to note that the bound on information transfer is precisely what is needed to prevent information loss.
+Because the throat capacity is of the same order as the Page-time interior entropy, we interpret this capacity as an upper bound for the extraction of information through the Hayden–Preskill protocol [18].
+Thus, a traversable wormhole might be used to support a quantum channel capable of transmitting all interior black hole degrees of freedom that remain entangled with the Hawking radiation, offering a way to extract all the information trapped behind the horizon.
+It is interesting to note that the bound on information transfer is consistent with unitarity constraints.
 It would be fruitful to explore this rather speculative topic on mathematical grounds.